# Model-Based Reward Shaping for Adversarial Inverse Reinforcement Learning in Stochastic Environments

### v1

@@ Abstract
In this paper, we aim to tackle the limitation of the Adversarial Inverse Reinforcement Learning (AIRL) method in stochastic environments where theoretical results cannot hold and performance is degraded. To address this issue, we propose a novel method which infuses the dynamics information into the reward shaping with the theoretical guarantee for the induced optimal policy in the stochastic environments. Incorporating our novel model-enhanced rewards, we present a novel Model-Enhanced AIRL framework, which integrates transition model estimation directly into reward shaping. Furthermore, we provide a comprehensive theoretical analysis of the reward error bound and performance difference bound for our method. The experimental results in MuJoCo benchmarks show that our method can achieve superior performance in stochastic environments and competitive performance in deterministic environments, with significant improvement in sample efficiency, compared to existing baselines.
@@ 1 Introduction
Reinforcement learning (RL) has achieved considerable success across various domains, including board game (Schrittwieser et al., 2020), MOBA game (Berner et al., 2019), time-delayed system (Wu et al., ; Wu et al., 2024), and cyber-physical systems (Wang et al., 2023a, b, c; Zhan et al., 2023). Despite these advances, RL highly depends on the quality of reward function design which demands expertise, intensive labour, and a great amount of time (Russell, 1998). To address this, imitation learning (IL) methods, such as Behavior Cloning (BC) (Torabi et al., 2018a) and Inverse Reinforcement Learning (IRL) (Arora & Doshi, 2021), leverage human or expert demonstrations to bypass the need for explicit reward functions. These methods aim to learn from the demonstrations to eventually match the distribution of expert behavior, and have shown great promise in applications like autonomous driving (Codevilla et al., 2018; Sun et al., 2018), legged locomotion (Peng et al., 2020; Ratliff et al., 2009), and planning tasks (Choudhury et al., 2018; Yin et al., 2022).
@@
The notable approaches within IRL are Adversarial Imitation Learning (AIL) methods that build upon maximum entropy framework (Ziebart et al., 2008). These adversarial methods frame imitation learning as a maximum likelihood estimation problem on trajectory distributions, converting the distribution into a Boltzmann distribution parameterized by rewards under deterministic environment settings (Wu et al., 2024). This closely mirrors the distribution approximation found in generative models (Finn et al., 2016a; Swamy et al., 2021). Thus, model-free AIL approaches often follow generative model structures, such as GANs (Ho & Ermon, 2016; Fu et al., 2017) or diffusion models (Reuss et al., 2023), and require extensive sampling for distribution matching gradient updates (Orsini et al., 2021) conducting. Model-based IRL and AIL frameworks have also emerged, where model-based framework is designed to provide estimation for gradient and planning, leading to innovative combinations such as gradient-based IRL with model predictive control (MPC) (Das et al., 2021) and end-to-end differentiable IRL frameworks for complex robotics tasks (Baram et al., 2016, 2017; Sun et al., 2021; Rafailov et al., 2021). However, these approaches primarily address deterministic settings and struggle when applied to stochastic environments.
@@
The only learning "deterministic" reward techniques among the existing AIL methods face significant performance degeneration in stochastic environments, leading to risk-seeking behavior and increased data requirements (Ziebart et al., 2010). For example, an agent trained under the deterministic Markov Decision Process (MDP) might aim to imitate expert behavior by seeking high rewards, yet fail to account for the low probability of some transitions in stochastic MDP settings. This happens because, in stochastic environments, the assumption that trajectory distributions are aligned with a Boltzmann distribution solely parameterized by deterministic rewards no longer holds. Instead, the dynamics information must also be incorporated into the formulation. There are two possible solutions. One is massive sampling to cover all possible outcomes, which is computationally expensive in large state action spaces (Devlin & Kudenko, 2011; Gupta et al., 2022). The other is estimating the dynamics information and integrating it into the reward design, making the reward "stochastic". Traditional reward design is usually based on state only $R(s_{t})$ (Torabi et al., 2018b), state-action pair $R(s_{t},a_{t})$ (Blondé & Kalousis, 2019), or transition tuple $R(s_{t},a_{t},s_{t+1})$ (Fu et al., 2017), where the information inputted can be thought as a deterministic sample piece under the stochastic setting. The challenge in stochastic environments calls for a different perspective of rewards – stochastic rewards absorbing the transition information.
Inspired by this idea, we propose a novel model-enhanced AIRL framework with a specifically tailored model-enhanced reward design approach to elevate performance in stochastic environments while remaining competitive in deterministic settings. In contrast to existing methods, our approach leverages the predictive power of the estimated transition model to shape rewards, represented as $\hat{R}(s_{t},a_{t},\hat{\mathcal{T}})$. This also enables us to generate synthetic trajectories to help guide policy optimization and reduce dependency on costly real-world interactions. As part of our analysis, we provide a theoretical guarantee on the optimal behavior for policies induced by our reward shaping and derive a bound on the performance gap with respect to the transition model errors. Empirically, we demonstrate that this integration significantly enhances sample complexity and policy performance in both settings, providing a comprehensive solution to the limitations of existing AIL methods in uncertain environments.
Contributions of this work include:
  • A novel reward shaping method with model estimation under the stochastic MDP setting, which provides the optimal policy invariance guarantee.
@@
  • A novel model-based adversarial IRL framework that seamlessly incorporates transition model training, adversarial reward learning with model estimation and forward model-based RL process, enhancing performance in stochastic environments, and sample efficiency.
@@
  • Theoretical analysis on reward learning with model estimation under the adversarial framework and performance difference under transition model learning errors.
@@
  • Empirical validation that demonstrates our approach’s performance improvements in stochastic environments and comparable performance in deterministic environments as well as significant improvement in sample efficiency.
In Sec. 2, we introduce related works in AIL and reward shaping. In Sec. 3, we provide the necessary preliminaries for MDP and IRL. In Sec. 4, we present our model-enhanced reward shaping method and corresponding theoretical guarantee. In Sec. 5, we present our Model-Enhanced AIRL framework design together with theoretical analysis on reward error bound and performance difference bound. In Sec. 6, we show the experimental results in Mujoco for various benchmarks. Sec. 7 concludes the paper.
2 Related Works
Adversarial Imitation Learning.
Margin optimization based IRL methods (Ng et al., 2000; Abbeel & Ng, 2004; Ratliff et al., 2006) aim to learn reward functions that explain expert behavior better than other policies by a margin. Bayesian approaches were introduced with different prior assumptions on reward distributions, such as Boltzmann distributions (Ramachandran & Amir, 2007; Choi & Kim, 2011; Chan & van der Schaar, 2021) or Gaussian Processes (Levine et al., 2011). Other statistical learning methods include multi-class classification (Klein et al., 2012; Brown et al., 2019) and regression trees (Levine et al., 2010). The entropy optimization approach has seen significant development. To avoid biases from maximum margin methods, the maximum entropy principle (Shore & Johnson, 1980) is used to infer distributions over trajectories parameterized by reward weights. Ziebart et al. (2008, 2010) proposed a Lagrangian dual framework to cast the reward learning into a maximum likelihood problem with linear-weighted feature-based reward representation. Wulfmeier et al. (2015) extended the framework to nonlinear reward representations, and Finn et al. (2016b) combines importance sampling techniques to enable model-free estimation. Inspired by GANs, adversarial methods were introduced for policy and reward learning in IRL (Ho & Ermon, 2016; Fu et al., 2017; Torabi et al., 2018b). However, these methods typically work with Maximum Entropy (ME) formulation yet suffer from sample inefficiency. Although there have been efforts to combine adversarial methods with off-policy RL agents to improve sample efficiency (Kostrikov et al., 2018; Blondé & Kalousis, 2019; Blondé et al., 2022), few extend it to the model-based setting which might further the improvement, and none of these approaches addresses the rewards learning in stochastic MDP settings.
Rewards Shaping.
Reward shaping (Dorigo & Colombetti, 1994; Randløv & Alstrøm, 1998) is a technique that enhances the original reward signal by adding additional domain information, making it easier for the agent to learn optimal behavior. This can be defines as $\hat{R}=R+F$, where $F$ is the shaping function and $\hat{R}$ is the shaped reward function. Potential-based reward shaping (PBRS) (Ng et al., 2000) builds the potential function on states, $F(s,a,s^{\prime})=\phi(s^{\prime})-\phi(s)$, while ensuring the policy invariance property, which refers to inducing the same optimal behavior under different rewards $R$ and $\hat{R}$. Nonetheless, there also exist other variants on the inputs of the potential functions such as state-action (Wiewiora et al., 2003), state-time (Devlin & Kudenko, 2012), and value function (Harutyunyan et al., 2015) as potential function input. There are also some latest attempts of reward shaping without utilization of domain knowledge potential function to solve exploration under sparse rewards (Hu et al., 2020; Devidze et al., 2022; Gupta et al., 2022; Skalse et al., 2023).
MBIRL.
Model-Based RL (MBRL) has emerged as a promising direction for improving sample efficiency and generalization (Janner et al., 2019; Yu et al., 2020). MBRL combines various learned dynamics neural network structures with planning (Hansen et al., 2022; Sikchi et al., 2022).This framework has been successfully extended to vision-based control tasks (Hafner et al., 2019; Zhan et al., 2023). Integrating IRL with MBRL has also shown success. For example, Das et al. (2021) and Herman et al. (2016) presented a gradient-based IRL approach using different policy optimization methods with dynamic models for linear-weighted features reward learning. In Das et al. (2021), the dynamic model is used to pass forward/backward the gradient in order to update the IRL and policy optimization modules. Similarly, end-to-end differentiable adversarial IRL frameworks to various state spaces have also been explored (Baram et al., 2016, 2017; Sun et al., 2021; Rafailov et al., 2021), where dynamic model serves a similar role. Despite these advancements, existing methods rarely address the specific challenges posed by stochastic environments, which limit reward learning performance. To our knowledge, this is the first study that provide a theoretical analysis on the performance difference with learned dynamic model for the adversarial IRL problem under stochastic MDP.
@@ 3 Preliminaries
RL is usually formulated as a Markov Decision Process (MDP) $\mathcal{M}$ (Puterman, 2014) denoted as a tuple $\langle\mathcal{S},\mathcal{A},\mathcal{T},\gamma,R,\rho_{0}\rangle$. $\rho_{0}$ is the initial distribution of the state. $s\in\mathcal{S},a\in\mathcal{A}$ stands for the state and action space respectively. $\mathcal{T}$ stands for the transition dynamic such that $\mathcal{T}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$. $\gamma\in(0,1)$ is the discounted factor, $R$ stands for reward function such that $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ and $\|R\|_{\infty}\leq R_{\max}$. The discounted visitation distribution of trajectory $\tau$ with policy $\pi$ is given by:
$p(\tau)=\rho_{0}\prod_{t=0}^{T-1}\gamma^{t}\mathcal{T}(s_{t+1}|s_{t},a_{t})\pi(a_{t}|s_{t}).$ (1)
The objective function of RL is $\max\mathbb{E}_{\tau\sim p(\tau)}\left[\sum_{t=0}^{T}\gamma^{t}R(\tau)-H(\pi)\right]$, where $H$ is the log likelihood of the policy. We introduce Soft Value Iteration for bellmen update (Haarnoja et al., 2018), where $Q^{soft}$ and $V^{soft}$ denotes the soft Q function and Value function respectively:
$V^{soft}(s_{t})$ $=log\sum_{a_{t}\in\mathcal{A}}\exp{Q^{soft}(s_{t},a_{t})}da_{t},$ (2)
$Q^{soft}(s_{t},a_{t})$ $=R(s_{t},a_{t})+\gamma\mathbb{E}_{\mathcal{T}}\left[V^{soft}(s_{t+1})|s_{t},a_{t}\right],$ (3)
$\pi(a_{t}|s_{t})$ $=\exp{(Q^{soft}(s_{t},a_{t})-V^{soft}(s_{t}))},$ (4)
where the soft Advantage function is defined as $A^{soft}(s_{t},a_{t})=Q^{soft}(s_{t},a_{t})-V^{soft}(s_{t})$.
@@
In IRL setting, we usually consider the MDP without reward as $\mathcal{M}^{\prime}$ where $R$ is also unknown. We denote the data buffer $\mathcal{D}_{exp}$ which collects trajectories from an expert policy $\pi^{E}$. We consider a reward function $R_{\theta}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, where $\theta$ is the reward parameter. An IRL problem can be defined as a pair $\mathcal{B}=(\mathcal{M}^{\prime},\pi^{E})$. A reward function $R_{\theta}$ is feasible for $\mathcal{B}$ if $\pi^{E}$ is an optimal policy for the MDP $\mathcal{M}^{\prime}\cup R_{\theta}$, and we denote the set of feasible rewards as $\mathcal{R}_{\mathcal{B}}$. Using maximize likelihood estimation framework, we can formulate the IRL as the following maximum causal entropy problem:
$\arg\max_{\theta}\mathbb{E}_{\tau\sim\mathcal{D}_{exp}}\log p_{\theta}(\tau),$ (5)
where $Q_{R_{\theta}}^{soft}$ and $V^{soft}_{R_{\theta}}$ are based on $R_{\theta}$ and $p_{\theta}(\tau)\propto\rho_{0}\prod_{t=0}^{T-1}\mathcal{T}(s_{t+1}|s_{t},a_{t})\exp(Q^{soft}_{\hat{R}_{\theta}}(s_{t},a_{t})-V^{soft}_{R_{\theta}}(s_{t}))$ (Ziebart et al., 2010). Under deterministic MDP, the above problem can be simplified as maximum entropy, where $p_{\theta}(\tau)\propto\frac{1}{Z_{\theta}}\exp{\sum_{t=0}^{T-1}R_{\theta}(s_{t},a_{t})}$ and $Z_{\theta}$ is the temperature factor of the Boltzmann Distribution (Ziebart et al., 2008).
4 Model Estimation in reward shaping
Table 1: We summarize the different reward formulations and their dynamic properties in this table. Components refer to the input pair that the reward functions take. Reward Shaping indicates whether there is the extra physical potential information involved where X means no reward shaping used. Dynamics information shows whether transitions are involved in the reward function.
Methods Components Reward Shaping Dynamics Information
AIRL (Fu et al., 2017) $s_{t},a_{t},s_{t+1}$ $R(s_{t},a_{t})+\gamma\phi(s_{t+1})-\phi(s_{t})$ single sample
AIRL(State Only) $s_{t}$ $R(s_{t})+\textrm{constant}$ X
DAC (Kostrikov et al., 2018) $s_{t},a_{t}$ X X
SAM (Blondé & Kalousis, 2019) $s_{t},a_{t}$ X X
SQIL (Reddy et al., 2019) $s_{t},a_{t}$ binary X
GAIfO (Torabi et al., 2018b) $s_{t},s_{t+1}$ X single sample
@@ 4 Model Estimation in reward shaping
In this section, we illustrate the advantages of involving transition dynamics into the reward shaping, especially in stochastic MDP settings. Most of literature work has various formulations and definitions (Table 1), but few considers transition dynamic information in the reward shaping. Defining rewards solely based on states, $R^{s}(s_{t})$, offers limited utility in environments where actions are critical. Even though the state-action pair-based rewards $R^{sa}(s_{t},a_{t})$ can capture the missing information of the taken action, it fails to consider any future information, the successive state $s_{t+1}$. Transition tuple-based rewards $R^{tuple}(s_{t},a_{t},s_{t+1})$ incorporate the dynamics information in a sampling-based way, which requires abundant data to learn the underlying relationship of two consecutive states, potentially raising the sample efficiency issue in the stochastic environment with the huge state space. To address this issue, we propose dynamics-based rewards shaping $\hat{R}(s_{t},a_{t},\mathcal{T})$, which explicitly infuse the dynamics information $\mathcal{T}$ on the potential function, thus significantly improving the sample-efficiency. Specifically, our rewards shaping is defined as
@@
where $\phi$ is a state-only potential function, $\mathcal{T}$ is the dynamics. Another insight of the above reward shaping is to resemble the advantage function with the soft value function as the potential function, which we will elaborate in Sec. 5.1. With the given reward shaping $\hat{R}$, it is crucial to show that it induces the same optimal behaviour as the ground-true reward $R$. We formally define this policy invariance property as follows.
@@
(Memarian et al., 2021) Let $R$ and $\hat{R}$ be two reward functions. We say they induce the same soft optimal policy under transition dynamics $\mathcal{T}$ if, for all states $s\in\mathcal{S}$ and actions $a\in\mathcal{A}$:
$A_{R}^{soft}(s_{t},a_{t})=A_{\hat{R}}^{soft}(s_{t},a_{t}).$ (7)
With the above definition, we can transfer the proof of policy invariant property of our designed reward shaping (Eq. 6) to showing the equivalence of soft advantage functions, which is proved in the following theorem. The detailed proof can be found in A.1.
Theorem 4.2 (Reward Invariance).
@@
Thm. 4.2 implies that the optimal policy induced from our model-enhanced rewards shaping $\hat{R}$ (Eq. 6) is equivalent to the optimal policy trained by the ground-truth reward function $R$ under the soft Value Iteration fashion.
5 Model Enhanced Adversarial IRL
In this section, we first elaborate on the adversarial formulation of our reward shaping (Eq. 8) and present the theoretical insight (Proposition 5.1) of the equivalence between cross-entropy training loss of adversarial reward shaping formulation and maximum log-likelihood loss of original maximum causal entropy IRL problem. Then, in the Sec. 5.2, we showcase our practical algorithm framework with trajectory generation and transition model learning in the loop, as shown in  Fig. 1. Furthermore, we theoretically investigate the reward function bound (Thm. 5.3) and performance difference bound (Thm. 5.4) under the transition model learning error.

### v2

# Enhancing Inverse Reinforcement Learning through Encoding Dynamic Information in Reward Shaping

@@ Abstract
Adversarial-based inverse reinforcement learning (IRL) has shown promising results using reward shaping under deterministic settings. However, it struggles in stochastic environments where existing theoretical results no longer apply, leading to degraded performance. To address this issue, we propose a novel maximum causal entropy based off-policy IRL method with transition-aware reward shaping framework. Our method integrates transition model estimation directly to learn stochastic-invariant rewards. We conduct a thorough theoretical analysis, establishing bounds on reward error and performance differences to validate the effectiveness of our method. The experimental results in continuous locomotion tasks (MuJoCo) show that our method can achieve superior performance in stochastic environments and competitive performance in deterministic environments, with significant improvement in sample efficiency, compared to existing baselines. Additionally, we extend our framework to high-dimensional vision-based tasks, where our method shows promising results on multiple stochastic Atari games. These results demonstrate that embedding transition awareness into reward learning is critical for robust IRL in realistic stochastic settings.
keywords:
Inverse Reinforcement Learning; Reinforcement Learning; Stochastic MDP
@@ 1 Introduction
Reinforcement learning (RL) has achieved considerable success across various domains, including board game (schrittwieser2020mastering), MOBA game (berner2019dota), time-delayed system (wuboosting; wu2024variational), and cyber-physical systems (wang2023empowering; wang2023joint; wang2023enforcing; zhan2024state). Despite these advances, RL highly depends on the quality of reward function design which demands expertise, intensive labour, and a great amount of time (russell1998learning). To address this, imitation learning (IL) methods, such as Behavior Cloning (BC) (torabi2018behavioral) and Inverse Reinforcement Learning (IRL) (arora2021survey), leverage human or expert demonstrations to bypass the need for explicit reward functions. These methods aim to learn from the demonstrations to eventually match the distribution of expert behavior, and have shown great promise in applications like autonomous driving (codevilla2018end; sun2018fast), legged locomotion (peng2020learning; ratliff2009learning), and planning tasks (choudhury2018data; yin2022planning).
@@
The notable approaches within IRL are Generative Imitation Learning (GIL) methods that build upon maximum entropy framework (ziebart2008maximum). These methods frame imitation learning as a maximum likelihood estimation problem on trajectory distributions, converting the distribution into a Boltzmann distribution parameterized by rewards under deterministic environment settings (wu2024variational). This closely mirrors the distribution approximation found in generative models (finn2016connection; swamy2021moments). Thus, model-free deep IL approaches often follow generative model structures, such as GANs (ho2016generative; fu2017learning) or diffusion models (reuss2023goal; wu2024diffusing), and require extensive sampling for distribution/score function matching in on-policy fashion (orsini2021matters). Model-based IL frameworks have also emerged, where model-based framework is designed to provide estimation for gradient and planning, leading to innovative combinations such as gradient-based IRL with model predictive control (MPC) (das2021model) and end-to-end differentiable IRL frameworks for complex robotics tasks (baram2016model; baram2017end; sun2021adversarial; rafailov2021visual). However, these approaches primarily address deterministic settings and struggle when applied to stochastic environments.
@@
Existing IRL methods, rooted in their maximum entropy nature, mostly exclusively focus on learning "deterministic" reward techniques. These methods, however, face significant performance degeneration in stochastic environments, leading to risk-seeking behavior and increased data requirements (ziebart2010modeling). For example, an agent trained under the deterministic Markov Decision Process (MDP) might aim to imitate expert behavior by seeking high rewards, yet fail to account for the low probability of some transitions in stochastic MDP settings. This happens because, in stochastic environments, the assumption of maximum entropy no longer holds. The trajectory distributions are not aligned with a Boltzmann distribution solely parameterized by deterministic rewards. In this case, the dynamic information must also be incorporated into the formulation. There are two likely solutions. One is massive sampling to cover all possible outcomes, which is computationally expensive in large state action spaces (devlin2011theoretical; gupta2022unpacking). The other is changing from maximum entropy framework to maximum causal entropy framework, estimating the dynamics information, and integrating it into the reward design, making the reward "stochastic". Traditional reward design is usually based on state only $R(s_{t})$ (torabi2018generative), state-action pair $R(s_{t},a_{t})$ (blonde2019sample), or transition tuple $R(s_{t},a_{t},s_{t+1})$ (fu2017learning), where the information inputted can be thought as a deterministic sample piece under the stochastic setting. The challenge in stochastic environments calls for a different perspective of rewards – stochastic rewards absorbing the transition information. More detailed literature survey has been has been investigated in Appendix A.
[…]
@@
  • A novel model-based off-policy IRL framework rooted in maximum causal entropy theory that seamlessly incorporates transition model training, adversarial reward learning with model estimation and forward model-based RL process, enhancing performance in stochastic environments, and sample efficiency.
@@
  • Theoretical analysis on reward learning and performance difference under transition model learning errors within our framework.
@@
  • Empirical validation showing our approach’s performance improvements in stochastic environments as well as significant gains in sample efficiency and comparable performance in deterministic environments.
We begin by introducing the necessary preliminaries on Markov Decision Processes (MDPs) and inverse reinforcement learning (IRL). We then present our model-enhanced reward shaping method, along with its theoretical guarantee. Next, we describe the full model-enhanced IRL framework, including its derivation from the maximum causal entropy objective, and provide theoretical analysis on the reward error bound and performance difference bound. Experimental results are reported on a range of MuJoCo and Atari benchmarks, demonstrating the effectiveness of our approach. We conclude with a discussion of related work and a summary of our findings.
@@ 2 Preliminaries
RL usually considers a Markov Decision Process (MDP) $\mathcal{M}$ (puterman2014markov) denoted as a tuple $\langle\mathcal{S},\mathcal{A},\mathcal{T},\gamma,R,\rho_{0}\rangle$. $\rho_{0}$ is the initial distribution of the state. $s\in\mathcal{S},a\in\mathcal{A}$ stands for the state and action space respectively. $\mathcal{T}$ stands for the transition dynamic such that $\mathcal{T}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$. $\gamma\in(0,1)$ is the discounted factor, $R$ stands for reward function such that $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ and $\|R\|_{\infty}\leq R_{\max}$. The undiscounted visitation distribution of trajectory $\tau$ with policy $\pi$ is given by $p(\tau)=\rho_{0}\prod_{t=0}^{T-1}\mathcal{T}(s_{t+1}|s_{t},a_{t})\pi(a_{t}|s_{t}).$ The objective function of RL is defined as $\max\mathbb{E}_{\tau\sim p(\tau)}\left[\sum_{t=0}^{T}\gamma^{t}R(\tau)-H(\pi)\right]$, where $H$ is the log likelihood of the policy. We introduce Soft Value Iteration for bellmen update (haarnoja2018soft), where $Q^{soft}$ and $V^{soft}$ denotes the soft Q function and Value function respectively:
$V^{soft}(s_{t})=log\sum_{a_{t}\in\mathcal{A}}\exp{Q^{soft}(s_{t},a_{t})},$ $\quad Q^{soft}(s_{t},a_{t})=R(s_{t},a_{t})+\gamma\mathbb{E}_{\mathcal{T}}\left[V^{soft}(s_{t+1})|s_{t},a_{t}\right],$
$\pi(a_{t}|s_{t})=\exp{(Q^{soft}(s_{t},a_{t})-V^{soft}(s_{t}))},$ $\quad A^{soft}(s_{t},a_{t})=Q^{soft}(s_{t},a_{t})-V^{soft}(s_{t})$
@@
In the IRL setting, we usually consider the MDP without reward as $\mathcal{M}^{\prime}$ where $R$ is also unknown. We denote the data buffer $\mathcal{D}_{exp}$ which collects trajectories from an expert policy $\pi^{E}$. We consider a reward function $R_{\theta}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, where $\theta$ is the reward parameter. An IRL problem can be defined as a pair $\mathcal{B}=(\mathcal{M}^{\prime},\pi^{E})$. A reward function $R_{\theta}$ is feasible for $\mathcal{B}$ if $\pi^{E}$ is an optimal policy for the MDP $\mathcal{M}^{\prime}\cup R_{\theta}$, and we denote the set of feasible rewards as $\mathcal{R}_{\mathcal{B}}$. Using the maximize likelihood estimation framework, we can formulate the IRL as the following maximum causal entropy (MCE) problem, $\arg\max_{\theta}\mathbb{E}_{\tau\sim\mathcal{D}_{exp}}\log p_{\theta}(\tau)$, where $Q_{R_{\theta}}^{soft}$ and $V^{soft}_{R_{\theta}}$ are based on $R_{\theta}$ and $p_{\theta}(\tau)\propto\rho_{0}\prod_{t=0}^{T-1}\mathcal{T}(s_{t+1}|s_{t},a_{t})\exp(Q^{soft}_{R_{\theta}}(s_{t},a_{t})-V^{soft}_{R_{\theta}}(s_{t}))$ (ziebart2010modeling). Under deterministic MDP, the above problem can be simplified as maximum entropy (ME) problem, where $p_{\theta}(\tau)\propto\frac{1}{Z_{\theta}}\exp{\sum_{t=0}^{T-1}R_{\theta}(s_{t},a_{t})}$ and $Z_{\theta}$ is the temperature factor of the Boltzmann Distribution (ziebart2008maximum).
3 Transition-Aware Reward Shaping
Table 1: This table summarizes different reward formulations and their dynamic properties. Components refer to the input pairs for the reward functions. Reward Shaping indicates whether additional physical potential information is included (X means none). Dynamics specifies if transitions are considered in the reward function.
Methods Components Reward Shaping Dynamic Information
AIRL(State Only) (fu2017learning) $s_{t}$ $R(s_{t})+\textrm{constant}$ X
DAC (kostrikov2018discriminator) $s_{t},a_{t}$ X X
SAM (blonde2019sample) $s_{t},a_{t}$ X X
SQIL (reddy2019sqil) $s_{t},a_{t}$ binary X
GAIfO (torabi2018generative) $s_{t},s_{t+1}$ X single sample
@@ 3 Transition-Aware Reward Shaping
In this section, we illustrate the advantages of involving transition dynamics into the reward shaping, especially in stochastic MDP settings. Existing work has proposed various formulations and definitions (Table 1), but few considers transition dynamic information in the reward shaping. Defining rewards solely based on states, $R^{s}(s_{t})$, offers limited utility in environments where actions are critical. Even though the state-action pair-based rewards $R^{sa}(s_{t},a_{t})$ can capture the missing information of the action taken, it fails to consider future information, the successive state $s_{t+1}$. Transition tuple-based rewards $R^{tuple}(s_{t},a_{t},s_{t+1})$ incorporate dynamic information in a sampling-based way, which requires abundant data to learn the underlying relationship of two consecutive states, potentially raising the sample efficiency issue in the stochastic environment with the huge state space. To address this issue, we propose transition-aware reward shaping $\hat{R}(s_{t},a_{t},\mathcal{T})$, which explicitly infuses dynamic information $\mathcal{T}$ on the potential function, thus significantly improving sample efficiency. Specifically, our reward shaping is defined as
@@
where $\phi$ is a state-only potential function, $\mathcal{T}$ is the dynamics. Another insight of the above reward shaping is to resemble the advantage function with the soft value function as the optimal potential function, which we will elaborate on next section. With the given reward shaping $\hat{R}$, it is crucial to show that it induces the same optimal behavior as the original reward $R$. We formally define this policy invariance property as follows.
@@
(memarian2021self) Let $R$ and $\hat{R}$ be two reward functions. We say they induce the same soft optimal policy under transition dynamics $\mathcal{T}$ if, for all states $s\in\mathcal{S}$ and actions $a\in\mathcal{A}$$A_{R}^{soft}(s_{t},a_{t})=A_{\hat{R}}^{soft}(s_{t},a_{t}).$
With the above definition, we can transfer the proof of policy invariant property of our designed reward shaping (Equation 1) to showing the equivalence of soft advantage functions, which is proved in the following theorem. The detailed proof can be found in Theorem B.1.
Theorem 3.2 (Policy Invariance).
@@
Theorem 3.2 implies that the optimal policy induced from our transition-aware reward shaping $\hat{R}$ (Equation 1) is equivalent to the optimal policy trained by the ground-truth reward function $R$ under the soft value iteration fashion.
4 Model Enhanced IRL